OpenStudy (anonymous):

can some one help me understand this question i don't get how to do it. Find the solution, given the replacement set. 5x - 2y = 6 {(-2, -8), (0, -3), (1, -1), (2, 4), (4, 7)} A. {(-2, -8), (2, 4), (4, 7)} B. {(-2, -8), (0, -3), (4, 7)} C. {(-2, -8), (0, -3), (1, -1)} D. {(1, -1), (2, 4), (4, 7)}

OpenStudy (cwrw238):

5(-2) - 2(-8) = -10+16 = 6 5(0) - 2(-3) = 6 5(4) - 2(7) = 6 B

OpenStudy (anonymous):

xœ-ŒA €0¿’cW,ÒkA¿ªÝ‚P=Tý¿Kõ’ÀIÖ‚Í‘aA‰ÄŒšŽ5'‘Ä

OpenStudy (cwrw238):

yeah

OpenStudy (turingtest):

the question is asking you to see which of the solutions {(-2, -8), (0, -3), (1, -1), (2, 4), (4, 7)} satisfy the equation 5x - 2y = 6 so cwrw's answer is correct

OpenStudy (anonymous):

In fact, all of the answers satisfy the equation